# Efficient Sampling for Pauli Measurement-Based Shadow Tomography in Direct Fidelity Estimation

### v5

@@ Abstract
A constant number of random Clifford measurements allows the classical shadow protocol to perform direct fidelity estimation (DFE) with high precision. However, estimating properties of an unknown quantum state is expected to be more feasible with random Pauli measurements than with random Clifford measurements in the near future. Inspired by the importance sampling technique applied to sampling Pauli measurements for DFE, we show that similar strategies can be derived from classical shadows. Specifically, we describe efficient methods using only local Pauli measurements to perform DFE with GHZ, W, and Dicke states, establishing tighter bounds (by factor of $14.22$ and $16$ for GHZ and W, respectively) on the number of measurements required for desired precision. These protocols are derived by adjusting the distribution of observables. Notably, they require no preprocessing steps other than the sampling algorithms.
@@ I Introduction
Advances in the preparation of many-body entangled quantum states have marked significant progress in the field of quantum information [1, 2, 3, 4]. A critical step in these experiments is verifying that the system’s state aligns with the desired one [5, 6, 7, 8]. Traditional methods for assessing quantum state fidelity include quantum state tomography (QST), which reconstructs the full density matrix of a quantum state, offering detailed information [9, 10, 11, 12]. However, QST is resource-intensive and scales poorly with system size, posing significant challenges for large-scale quantum systems [13]. Recent approaches have introduced various techniques aimed at simplifying the fidelity estimation process [14, 15, 16]. Among these, direct fidelity estimation (DFE) has gained attention for its potential to provide accurate fidelity measurements with fewer resources compared to full state tomography. Unlike full tomography, DFE focuses on directly estimating the fidelity without requiring a complete reconstruction of the quantum state, thus significantly reducing the measurement overhead. For some classes of states, efficient methods using Pauli measurement [14, 15, 17], some relying on importance sampling [18, 19], are known.
@@
Meanwhile, the classical shadow estimation technique is known for its pioneering approach to efficiently predicting numerous properties of unknown quantum states [20]. Specifically, it can estimate any quantity related to a quantum state $\rho$, expressed as $\text{tr}(\rho O)$ for some linear operator $O$. The fidelity between $\rho$ and a particular pure state is in this form, which implies that classical shadows can be used for DFE.
@@
A classical snapshot of $\rho$ is generated by performing a random unitary operation $U$ on $\rho$ followed by a computational basis measurement. The efficiency of classical shadow tomography is determined by the distribution from which $U$ is chosen. Theoretically, random Clifford measurements [21, 22, 23, 24] are preferable in the sense that a constant number of measurements suffices to estimate the fidelity with any pure target state. Detailed analyses of the performance of this measurement have been carried out in the context of DFE. Unfortunately, performing random Clifford measurements in a laboratory setting is currently impractical. More feasible alternatives are random Pauli measurements [24]. They are much more resource-efficient and need precise control over only single-qubit operations. The significant drawback of random Pauli measurements is that, in the worst case, exponentially many samples with respect to the locality of an observable are necessary for precise estimation [20]. However, for observables associated with the GHZ state and the W state, which have maximum locality, the number of measurements required is only polynomial in the number of qubits.
@@
In this work, we show that for some classes of target states, random Pauli measurements can be efficiently synthesized with classical shadows for DFE. Our protocols are designed to improve previous methods (based on characteristic functions) in the nonadaptive local Pauli measurement setting [14]. The central aspect is that we employ a target-specific distribution of Pauli measurements rather than the default uniform distribution. This is analogous to sampling measurements from the characteristic function $\chi_{\rho}(k)$, which depends on the target $\rho$. Accordingly, the shadow inversion step is adapted to align with the modified distribution of measurements. As a result, the bound on the number of measurements necessary to achieve a specified precision is tightened by a factor between 14.22 and 28.44, and between 16 and 32 for the GHZ state [25] and the W state [26, 27], respectively. The proposed estimator is more direct than Ref. [14] (as will be explained in the discussion), resulting in more accurate estimation with the same number of measurements. Moreover, it does not require any preprocessing steps or memory to store the distribution. The proposed method can be extended to other scenarios such as [28, 29].
@@ IV.1 Random $n$-qubit Clifford circuits
where $\mathbb{I}$ denotes the $2^{n}$-dimensional identity operator. A classical shadow of size independent of $n$ is sufficient to accurately estimate the fidelity with any pure target state [20, Theorem 1].
@@ IV.2 Tensor products of random single-qubit Clifford circuits
The worst case upper bound on the number of measurements to estimate (3) scales exponentially with the locality of $O$ [20, Proposition 3] or the stabilizer Rényi entropy [17, Theorem 1]. However, there exist much efficient methods for DFE with some well-conditioned states (as defined in Ref. [14]), despite the locality of $O$ being $n$.
@@ V.1 Fidelity with computational basis states
$\langle b_{1}|\hat{\rho}|b_{2}\rangle=\prod_{i=1}^{n}(3\langle b_{1i}|U_{i}^{\dagger}|\hat{b}_{i}\rangle\langle\hat{b}_{i}|U_{i}|b_{2i}\rangle-\delta_{b_{1i},b_{2i}}).$ (6)
@@ V.2 Fidelity with GHZ state
Input: Unknown state $\sigma$, error parameters $(\epsilon,\delta)$
Output: Estimate of the fidelity $F(\rho,\sigma)$, where $\rho$ is the $n$-qubit GHZ state
@@
    $\sigma^{\prime}\leftarrow U\sigma U^{\dagger}$
    $\hat{b}=\texttt{measureComputationalBasis}(\sigma^{\prime})$
@@
$N=\left\lceil\frac{9\log(2/\delta)}{8\epsilon^{2}}\right\rceil$
@@
The total number of measurements is about $9\log(2/\delta)/8\epsilon^{2}$, which tightens the bound (2) by a factor of $9128(1+\log_{2/\delta}2)/$ 9, a constant between $128/9$ and $256/9$ (closer to $128/9$ for small $\delta$). Alg. 1 provides a complete pseudocode for the GHZ state. It should be noted that Alg. 1 includes the Bell state verification described in Ref. [34] as a special case, where the measurement setting is chosen uniformly from $\{XX,-YY,ZZ\}$.
@@ V.3 Fidelity with W state
For the off-diagonal part, consider the term
From Table 1, we see that the term inside the square bracket in (V.3) is zero if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=0$, ±$\pm 1/4$ if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=1$ and $U_{i}=U_{j}$, and ±$\pm i/4$ if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=1$ and $U_{i}\neq U_{j}$. Meanwhile, $\text{tr}(\rho|e_{j}\rangle\langle e_{i}|)$ is the complex conjugate of (V.3), and the sample
$V_{o}=\left(\langle 0|U_{i}^{\dagger}|\hat{b}_{i}\rangle\langle\hat{b}_{i}|U_{i}|1\rangle\langle 1|U_{j}^{\dagger}|\hat{b}_{j}\rangle\langle\hat{b}_{j}|U_{j}|0\rangle+\langle 0|U_{j}^{\dagger}|\hat{b}_{j}\rangle\langle\hat{b}_{j}|U_{j}|1\rangle\langle 1|U_{i}^{\dagger}|\hat{b}_{i}\rangle\langle\hat{b}_{i}|U_{i}|0\rangle\right)\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}$ (13)
$o=\sum_{i<j}\sum_{\begin{subarray}{c}U_{i}=U_{j}\in T\setminus\{\mathbb{I}\},\\ U_{k\notin\{i,j\}}=\mathbb{I}\end{subarray}}\mathbb{E}_{\hat{b}}\left[\frac{V_{o}}{n}\right].$
where $\hat{o}\equiv(n-1)V_{o}$. Therefore, we have $0\leq\hat{d}\leq 1/n$ and $-(n-1)/2\leq\hat{o}\leq(n-1)/2$.
@@
Input: Unknown state $\sigma$, error parameters $(\epsilon,\delta)$
Output: Estimate of the fidelity $F(\rho,\sigma)$, where $\rho$ is the $n$-qubit W state
@@
    $\sigma^{\prime}\leftarrow U\sigma U^{\dagger}$
    $\hat{b}=\texttt{measureComputationalBasis}(\sigma^{\prime})$
@@
$N=\left\lceil\frac{\log(2/\delta)(n^{2}-n+1)^{2}}{2\epsilon^{2}n^{2}}\right\rceil$
@@
The total number of measurements is about $\log(2/\delta)n^{2}/2\epsilon^{2}$, which tightens the bound (2) by a factor of $16(1+\log_{2/\delta}2)$, a constant between $16$ and $32$ (closer to $16$ for small $\delta$). Alg. 2 provides a complete pseudocode for the W state.

### v6

@@ Abstract
A constant number of random Clifford measurements allows the classical shadow protocol to perform direct fidelity estimation (DFE) with high precision. However, estimating properties of an unknown quantum state is expected to be more feasible with random Pauli measurements than with random Clifford measurements in the near future. Inspired by the importance sampling technique applied to sampling Pauli measurements for DFE, we show that similar strategies can be derived from classical shadows. Specifically, we describe efficient methods using only local Pauli measurements to perform DFE with GHZ, W, and Dicke states, establishing tighter bounds (by factor of $14.22$ and $16$ for GHZ and W/Dicke, respectively) on the number of measurements required for desired precision. These protocols are derived by adjusting the distribution of observables. Notably, they require no preprocessing steps other than the sampling algorithms.
@@ I Introduction
Advances in the preparation of many-body entangled quantum states have marked significant progress in the field of quantum information [1, 2, 3, 4]. A critical step in these experiments is verifying that the system’s state aligns with the desired one [5, 6, 7, 8]. Traditional methods for assessing quantum state fidelity include quantum state tomography (QST), which reconstructs the full density matrix of a quantum state, offering detailed information [9, 10, 11, 12, 13, 14, 15, 16, 17]. However, QST is resource-intensive and scales poorly with system size, posing significant challenges for large-scale quantum systems [18]. Recent approaches have introduced various techniques aimed at simplifying the processes of verification [19, 20, 21, 22] and fidelity estimation [23, 24, 25]. Among these, direct fidelity estimation (DFE) has gained attention for its potential to provide accurate fidelity measurements with fewer resources compared to full state tomography. Unlike full tomography, DFE focuses on directly estimating the fidelity without requiring a complete reconstruction of the quantum state, thus significantly reducing the measurement overhead. For some classes of states, efficient methods using Pauli measurement [26, 23, 24, 27], some relying on importance sampling [28, 29], are known.
@@
Meanwhile, the classical shadow estimation technique is known for its pioneering approach to efficiently predicting numerous properties of unknown quantum states [30], and recent studies have begun to explore its variants [31, 32]. Specifically, it can estimate any quantity related to a quantum state $\rho$, expressed as $\text{tr}(\rho O)$ for some linear operator $O$. The fidelity between $\rho$ and a particular pure state is in this form, which implies that classical shadows can be used for DFE.
@@
A classical snapshot of $\rho$ is generated by performing a random unitary operation $U$ on $\rho$ followed by a computational basis measurement. The efficiency of classical shadow tomography is determined by the distribution from which $U$ is chosen. Theoretically, random Clifford measurements [33, 34, 35, 36] are preferable in the sense that a constant number of measurements suffices to estimate the fidelity with any pure target state. Detailed analyses of the performance of this measurement have been carried out in the context of DFE. Unfortunately, performing random Clifford measurements in a laboratory setting is currently impractical [37]. More feasible alternatives are random Pauli measurements [36]. They are much more resource-efficient and need precise control over only single-qubit operations. The significant drawback of random Pauli measurements is that, in the worst case, exponentially many samples with respect to the locality of an observable are necessary for precise estimation [30]. However, for observables associated with the GHZ state [38] and the W/Dicke state [39, 40, 41], which have maximum locality, the number of measurements required is only polynomial in the number of qubits. These entangled states play a central role as computational resources in various quantum protocols [42, 43, 44, 45, 46].
@@
In this work, we show that for some classes of target states, random Pauli measurements can be efficiently synthesized with classical shadows for DFE. Our protocols are designed to improve previous methods (based on characteristic functions) in the nonadaptive local Pauli measurement setting [23]. The central aspect is that we employ a target-specific distribution of Pauli measurements rather than the default uniform distribution. This is analogous to sampling measurements from the characteristic function $\chi_{\rho}(k)$, which depends on the target $\rho$. Accordingly, the shadow inversion step is adapted to align with the modified distribution of measurements. As a result, the bound on the number of measurements necessary to achieve a specified precision is tightened by a factor between 14.22 and 28.44, and between 16 and 32 for the GHZ state and the W/Dicke state, respectively. The proposed estimator is more direct than Ref. [23] (as will be explained in the discussion), resulting in more accurate estimation with the same number of measurements. Moreover, it does not require any preprocessing steps or memory to store the distribution. The proposed method is analyzed in a more general setting in Section VI, and it can also be extended to other scenarios as in [47, 48].
@@ IV.1 Random $n$-qubit Clifford circuits
where $\mathbb{I}$ denotes the $2^{n}$-dimensional identity operator. A classical shadow of size independent of $n$ is sufficient to accurately estimate the fidelity with any pure target state [30, Theorem 1].
@@ IV.2 Tensor products of random single-qubit Clifford circuits
The worst case upper bound on the number of measurements to estimate (3) scales exponentially with the locality of $O$ [30, Proposition 3] or the stabilizer Rényi entropy [27, Theorem 1]. However, there exist much efficient methods for DFE with some well-conditioned states (as defined in Ref. [23]), despite the locality of $O$ being $n$.
@@ V.1 Fidelity with computational basis states
$\langle b_{1}|\hat{\rho}|b_{2}\rangle=\prod_{i=1}^{n}(3\langle b_{1,i}|U_{i}^{\dagger}|\hat{b}_{i}\rangle\langle\hat{b}_{i}|U_{i}|b_{2,i}\rangle-\delta_{b_{1,i},b_{2,i}}).$ (6)
@@ V.2 Fidelity with GHZ state
Input: Unknown state $\rho$, error parameters $(\epsilon,\delta)$
Output: Estimate of the fidelity $F(\rho,\sigma)$, where $\sigma$ is the $n$-qubit GHZ state
@@
    $\rho^{\prime}\leftarrow U\rho U^{\dagger}$
    $\hat{b}=\texttt{measureComputationalBasis}(\rho^{\prime})$
@@
$N=\left\lceil\frac{9\log(2/\delta)}{8\epsilon^{2}}\right\rceil$ (12)
@@
The total number of measurements is about $9\log(2/\delta)/8\epsilon^{2}$, which tightens the bound (2) by a factor of $9128(1+\log_{2/\delta}2)/$ 9, a constant between $128/9$ and $256/9$ (closer to $128/9$ for small $\delta$). Alg. 1 provides a complete pseudocode for the GHZ state. It should be noted that the measurement settings for fidelity estimation in Alg. 1 also appear in some verification protocols [53, 54], but in our context these settings are further associated with a fidelity estimator defined by the specific measurement basis and outcome.
@@ V.3 Fidelity with W state
The off-diagonal estimator $\hat{o}$ can be found in Appendix A. We have $0\leq\hat{d}\leq 1/n$ and $-(n-1)/2\leq\hat{o}\leq(n-1)/2$.
@@
Input: Unknown state $\rho$, error parameters $(\epsilon,\delta)$
Output: Estimate of the fidelity $F(\rho,\sigma)$, where $\sigma$ is the $n$-qubit W state
@@
    $\rho^{\prime}\leftarrow U\rho U^{\dagger}$
    $\hat{b}=\texttt{measureComputationalBasis}(\rho^{\prime})$
@@
$N=\left\lceil\frac{\log(2/\delta)(n^{2}-n+1)^{2}}{2\epsilon^{2}n^{2}}\right\rceil$ (13)
@@
The total number of measurements is about $\log(2/\delta)n^{2}/2\epsilon^{2}$, which tightens the bound (2) by a factor of $16(1+\log_{2/\delta}2)$, a constant between $16$ and $32$ (closer to $16$ for small $\delta$). Alg. 2 provides a complete pseudocode for the W state. For relevant definitions, see Appendix A.
@@ V.4 Fidelity with Dicke state
We have $0\leq\hat{d}\leq\binom{n}{k}^{-1}$ and $-\binom{n}{k}^{-1}\cdot c_{l}\leq\hat{o}_{l}\leq\binom{n}{k}^{-1}\cdot c_{l}$, where the definition of $c_{l}$ and the off-diagonal estimator $\hat{o}_{l}$ can be found in Appendix B. Moreover, let
$S\equiv\frac{1}{2}+\sum_{l=\max(0,2k-n)}^{k-1}c_{l}.$
$=d+\sum_{l=\max(0,2k-n)}^{k-1}o_{l}$
$=d+o$
$=F(\rho,\sigma).$
We generate
samples of $\hat{F}$. From Hoeffding’s inequality,
Table 4: ALG. 3. DFE for $n$-qubit Dicke state with $k$ excitations.
Input: Unknown state $\rho$, error parameters $(\epsilon,\delta)$
Output: Estimate of the fidelity $F(\rho,\sigma)$, where $\sigma$ is the $n$-qubit Dicke state with $k$ excitations
Define $c_{l}$ as in (21).
$\texttt{sum}\leftarrow 0$
for    in range($N$):
   if $l=k$:
    $\hat{b}=\texttt{measureComputationalBasis}(\rho)$
[…]
   else
    $U\leftarrow\bigotimes_{m=1}^{n}U_{m}$
    $\rho^{\prime}\leftarrow U\rho U^{\dagger}$
    $\hat{b}=\texttt{measureComputationalBasis}(\rho^{\prime})$
            + |$|\{m\in\text{s}(\mathbf{j})-\text{s}(\mathbf{i})\mid\hat{b}_{m}=1\}|$
[…]
   end if
end for
return $\tilde{F}=\,\texttt{sum}/N+1/2\binom{n}{k}$
The total number of measurements is $O(n^{2k})$, where the asymptotics are with respect to $n$, while $k$, $\epsilon$, and $\delta$ remain constant. As a detailed comparison, it follows from [23] that estimating the fidelity with an $n$-qubit Dicke state $\sigma$ with $k$ excitations to additive error $\epsilon$ and failure probability $\leq\delta$ needs roughly
$\frac{8\log(4/\delta)}{\epsilon^{2}}\binom{n}{k}^{2}\approx\frac{8\log(4/\delta)}{\epsilon^{2}}\frac{n^{2k}}{(k!)^{2}}$ (14)
Pauli measurements. This follows from the fact that for every Pauli string $W\in\{I,X,Y,Z\}^{\otimes n}$,
$|\text{tr}(\sigma W)|=0\quad\text{or}\quad|\text{tr}(\sigma W)|\geq\binom{n}{k}^{-1}.$
Our protocol leverages Dicke state symmetries to reduce the sample space. For large $n$ (with fixed $k$) we have
$S\approx c_{0}=\binom{n}{2k}\binom{2k}{0}\binom{2k-1}{k-1}\approx\frac{n^{2k}}{2(k!)^{2}}.$
Consequently, the required number of samples becomes
$\frac{2\log(2/\delta)}{\epsilon^{2}}\frac{S^{2}}{\binom{n}{k}^{2}}\approx\frac{2\log(2/\delta)}{\epsilon^{2}}\frac{n^{4k}}{4(k!)^{4}}\frac{(k!)^{2}}{n^{2k}}=\frac{\log(2/\delta)}{2\epsilon^{2}}\frac{n^{2k}}{(k!)^{2}}.$ (15)
Equation (15) improves upon (14) by the factor $16(1+\log_{2/\delta}2)$, which recovers the W state gain when $k=1$. Alg. 3 provides a complete pseudocode for the Dicke state with $k$ excitations. For relevant definitions, see Appendix B.
VI A generic optimization rule
The protocols developed for GHZ, W, and Dicke states can be subsumed under a more general optimization framework. Consider a real-valued observable
$O=\sum_{b_{1},b_{2}\in\{0,1\}^{n}}c_{b_{1},b_{2}}|b_{1}\rangle\langle b_{2}|,\quad c_{b_{1},b_{2}}\in\mathbb{R}.$
For a fixed pair $(b_{1},b_{2})$ define
$\text{eq}(b_{1},b_{2})=\{i\mid b_{1,i}=b_{2,i}\},\quad\text{neq}(b_{1},b_{2})=\{i\mid b_{1,i}=1-b_{2,i}\}.$
To estimate the symmetric combination $\langle b_{1}|\rho|b_{2}\rangle+\langle b_{2}|\rho|b_{1}\rangle$, it suffices to perform a Pauli measurement that meets the following local basis rule:
  • qubits in $\text{eq}(b_{1},b_{2})$ are measured in the $Z$ basis.
  • qubits in $\text{neq}(b_{1},b_{2})$ are measured in the $X$ or $Y$ basis, with even number of $Y$ basis across $\text{neq}(b_{1},b_{2})$.
Let $\mathcal{P}^{(b_{1},b_{2})}$ be the set of Pauli strings obeying this rule. Because $O$ has real coefficients (so $c_{b_{1},b_{2}}=c_{b_{2},b_{1}}$ when $b_{1}\neq b_{2}$), the expectation value can be reconstructed solely from
$\bigcup_{\begin{subarray}{c}b_{1},b_{2}\in\{0,1\}^{n},\\ c_{b_{1},b_{2}}\neq 0\end{subarray}}\mathcal{P}^{(b_{1},b_{2})}.$ (16)
After restricting the sampling domain to (16) we reweight and renormalize the estimators and the sampling probabilities (which is itself a nontrivial process that depends on both the set (16) and the coefficients $c_{b_{1},b_{2}}$), replacing the original shadow estimation using all $3^{n}$ Pauli operators. An immediate corollary is that Pauli strings containing an odd number of $Y$ are never required when $O$ is real.
Whenever the target state factorizes as a tensor product of independent subsystems—such as a collection of Bell pairs used as essential resources in distributed quantum computing [55, 56, 57]—one incurs no loss of optimality by simply performing the optimal measurement protocol on each subsystem in parallel. Moreover, to estimate every subsystem fidelity to within an additive error $\epsilon$ with confidence at least $1-\delta$, the total number of measurements need only grow logarithmically with the number of subsystems.
VII Numerical simulations
Using the previous importance sampling technique, the average number of samples falls strictly below the bound (2) for many configurations. Furthermore, the bound (2) itself may be quite rough, as it involves two layers of estimation: one for importance sampling and another for characteristic function estimation. Thus, it is essential to experimentally demonstrate the improvement of our method over the importance sampling approach.
In practice, the precision of estimates in Algs 1 and 2 is not directly controllable, but the number of samples can be specified. Therefore, we compare Ref. [23] with Algs 1 and 2 as follows: First, Ref. [23] with parameters $(\epsilon,\delta)$ is used to estimate the fidelity. Next, Algs 1 and 2 are executed using the same number of samples as Ref. [23]. Finally, the mean squared error (MSE) of the estimates is evaluated as $\mathbb{E}\left[\left(\tilde{F}-F(\rho,\sigma)\right)^{2}\right]$, where $\tilde{F}$ is the output of our estimation protocol.
Input: Target state $\sigma$, fidelity $f$
Output: Random state $\rho$ such that $F(\rho,\sigma)=f$
$\rho$ = randomDensityMatrix()
$P\leftarrow I-\sigma$
$\rho\leftarrow P\rho P$
$\rho\,\texttt{*=}\,(1-f)/\text{tr}(\rho)$
return $\rho+f\sigma$
Figure 2: FIG. 2. MSE comparison for (a) GHZ state with $(\epsilon,\delta)=(0.1,0.1)$, (b) GHZ state with $(\epsilon,\delta)=(0.05,0.05)$, (c) W state with $(\epsilon,\delta)=(0.1,0.1)$, and (d) W state with $(\epsilon,\delta)=(0.05,0.05)$. Blue corresponds to the algorithm in Ref. [23], which is based on Pauli-Liouville coefficients [58], and orange to our approach. Error bars represent 95% confidence intervals.
The states were randomly generated with fidelities ranging from 0 to 1 in increments of 0.01 (i.e., $0,0.01,0.02,\ldots,1$), as described in Alg. 4. Each state is generated by first creating a random mixed state and projecting it onto the subspace orthogonal to the target state (GHZ or W). The output state with the desired fidelity is then obtained as a linear combination of the target state and the projected state. FIG. 2 shows the comparison of MSE for up to 7 qubits. Algorithms 1 and 2 significantly improve estimation accuracy over Ref. [23], although the gain in this direction is less pronounced than the reduction implied by (2) vs. (12) and (13). The simulations were performed using PennyLane’s classical simulation framework [59].
VIII Discussion
Practical measurements suffer from noise, and the robust shadow estimation protocol was proposed to specifically mitigate this source of error [58]. Its calibration procedure for uniformly random Pauli measurements requires a number of samples that grows exponentially with the locality of the observable. Since every observable considered in this work is $n$-local, a direct application of that result would incur prohibitive overhead. However, our protocol departs from uniform Pauli sampling by tailoring the measurement ensemble to each target observable, and we believe that one can similarly devise bespoke calibration schemes whose sample complexity does not scale exponentially in $n$. We therefore leave the development of such observable-specific calibration procedures as a key avenue for future work.
Acknowledgements.
Author Contributions
Appendix A Off-diagonal terms for W state
For the off-diagonal part, consider the term
From Table 1, we see that the term inside the square bracket in (A) is zero if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=0$, ±$\pm 1/4$ if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=1$ and $U_{i}=U_{j}$, and ±$\pm i/4$ if $\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}=1$ and $U_{i}\neq U_{j}$. Meanwhile, $\text{tr}(\rho|e_{j}\rangle\langle e_{i}|)$ is the complex conjugate of (A), and the sample
$V_{\text{W}}=\left(\langle 0|U_{i}^{\dagger}|\hat{b}_{i}\rangle\langle\hat{b}_{i}|U_{i}|1\rangle\langle 1|U_{j}^{\dagger}|\hat{b}_{j}\rangle\langle\hat{b}_{j}|U_{j}|0\rangle+\langle 0|U_{j}^{\dagger}|\hat{b}_{j}\rangle\langle\hat{b}_{j}|U_{j}|1\rangle\langle 1|U_{i}^{\dagger}|\hat{b}_{i}\rangle\langle\hat{b}_{i}|U_{i}|0\rangle\right)\delta_{\hat{b}_{[n]\setminus\{i,j\}},\mathbf{0}}$ (18)
$o=\sum_{i<j}\sum_{\begin{subarray}{c}U_{i}=U_{j}\in T\setminus\{\mathbb{I}\},\\ U_{k\notin\{i,j\}}=\mathbb{I}\end{subarray}}\mathbb{E}_{\hat{b}}\left[\frac{V_{\text{W}}}{n}\right].$
where $\hat{o}\equiv(n-1)V_{\text{W}}$. Therefore, we have $-(n-1)/2\leq\hat{o}\leq(n-1)/2$.
Appendix B Off-diagonal terms for Dicke state